The sum of two numbers is 12.if two times the added to three times the second is equal to 21,find the numbers. Answer by mananth(16949) 
2 x + 3 y = 21 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 1
-3 x -3 y = -36
2 x + 3 y = 21
Add the two equations
-1 x = -15
/ -1
x = 15
plug value of x in (1)
1 x + 1 y = 12
15 + y = 12
y = 12
y = -3
y = -3
x= 15 one number
y= -3 other number
m.ananth@hotmail.ca